# Coreset-Based Task Selection for Sample-Efficient Meta-Reinforcement Learning

### v1

@@ 1 Introduction
Figure 1: Comparison of coreset MAML-RL (this work) and MAML-RL on the hopper Mujoco environment (see appendix for details).
@@ 2.2.2 Coreset-based Task Selection
Definition 1 (Submodularity Nemhauser et al. (1978)).
@@ 4 Numerical Validation
Figure 2: Reward comparison of Algorithm 1 and the vanilla MAML (Finn et al., 2017) on RL cart pole tasks.
@@
Cart Pole: We evaluate Algorithm 1 in a deep meta-RL setting [^2] Code can be downloaded from: https://github.com/jd-anderson/Coreset-meta-RL.. In particular, we examine the cart pole environment (Towers et al., 2024), a classical control environment where physical properties of the system vary across tasks, including cart mass, pole mass, and pole length. The policy is parameterized by a multi-layer perceptron (MLP) architecture consisting of two hidden layers. Further details are provided in Appendix 6.9. Figure 2 shows the learning curves comparing Algorithm 1 with vanilla MAML (Finn et al., 2017), depicting both average rewards and standard deviations across 5 runs.
@@
Our results demonstrate that Algorithm 1 learns approximately 3×3$\times$3 × faster than the vanilla MAML algorithm, reaching higher reward values in fewer iterations. While our approach exhibits slightly higher variance (as indicated by the larger shaded regions representing standard deviation), it consistently outperforms the baseline in terms of learning speed. Notably, the task selection method reaches a reward of 150 around iteration $60$, whereas MAML takes approximately $200$ iterations to research the same reward. These empirical findings strongly support our theoretical analysis regarding sample complexity reduction (Corollary 1), validating that careful task selection significantly enhance sample-efficiency in meta-RL.
@@
LQR: We follow the setting proposed by Toso et al. (2024b) to validate our theoretical guarantees in the MAML-LQR setting. In particular, Figure 3 (left) shows the optimality gap across iterations. We implemented the MAML-LQR on three scenarios: the full task pool (40 tasks), a selected subset (10 tasks), and two ablation baselines - selected subset without weight assignment and randomly selected subset. Our results demonstrate the faster convergence on the weighted selected subset, while both unweighted selected subset and random subset achieves at most the same performance as the full task pool. Moreover, Figure 3 (right) depicts the optimality gap between the selected subset and full task pool with respect to sample size, confirming our theoretical results on sample complexity reduction for the MAML-LQR (Corollary 2).
@@
Figure 3: Optimality gap of Algorithm 1 in the MAML-LQR setting with respect to iterations.
@@ Acknowledgements
LT is funded by the Center for AI and Responsible Financial Innovation (CAIRFI) Fellowship and by the Columbia Presidential Fellowship. JA is partially funded by NSF grants ECCS 2144634 and 2231350.

### v2

@@ 1 Introduction
Figure 1: Comparison of coreset MAML-RL (this work) and MAML-RL on the walker2D Mujoco environment (see Section 4 for details).
@@ 2.2.2 Coreset-based Task Selection
Definition 1 (Submodularity (Nemhauser et al., 1978)).
@@ 4 Numerical Validation
Meta-Reinforcement Learning: We evaluate Algorithm 1 in a deep meta-RL setting. All experiments were implemented using PyTorch, OpenAI Gym (Towers et al., 2024) and Mujoco (Todorov et al., 2012). In particular, we examine the cartpole, hopper, and walker2D environments (Towers et al., 2024). These environments have physical properties of the system that vary across tasks, including the mass (cart pole, hopper, and walker2D), pole mass and length (cart pole) and friction coefficient (hopper and walker2D). The policy is parameterized by a multi-layer perceptron (MLP) architecture consisting of two hidden layers. For cart pole and hopper, we maintained an episodic task pool of 800 tasks with a selection ratio of 25%. For each iteration, we used a batch size of 10 tasks. For walker2D, which is a more complex task with higher dimension of observation and action space, we maintained an episodic task pool of 1600 tasks with a selection ratio of 20%. For each iteration, we used a batch size of 32 tasks. Figure 2 shows the learning curves comparing Algorithm 1 with vanilla MAML Finn et al. (2017), depicting both average rewards and standard deviations across 5 runs.
Figure 2: Reward comparison of Algorithm 1 and vanilla MAML (Finn et al., 2017) on Cart Pole (left), Hopper (middle), and Walker2D (right) tasks (Mujoco).
@@
Our results demonstrate that Algorithm 1 learns approximately 4×4$\times$4 × faster than the vanilla MAML algorithm, reaching higher reward values in fewer iterations. Our approach exhibits slightly higher variance (as indicated by the shaded regions representing standard deviation) in cart pole and walker2D and it consistently outperforms the baseline in terms of learning speed across the three tasks. Notably, in hopper, the task selection method reaches a reward of 120 around iteration $60$, whereas MAML takes approximately $200$ iterations to research a lower reward around $90$. In walker2D, the most complex task among the three tasks, our algorithm shows significant more than 4×4$\times$4 × speedup of reaching higher rewards during early training stage. These empirical findings strongly support our theoretical analysis regarding sample complexity reduction (Corollary 1), validating that careful task selection significantly enhance sample-efficiency in meta-RL. Moreover, all the experiments were running on an NVIDIA GeForce 3090 GPU with 90GB RAM. As we mentioned previously, selecting tasks via submodular maximization is efficient. In hopper, the average per-iteration running time was 3.72s for the coreset task selection algorithm and 3.35s for vanilla MAML.
@@
Linear Quadratic Regulator: We follow the setting proposed by (Toso et al., 2024b) to validate our theoretical guarantees in the MAML-LQR setting. In particular, Figure 3 (left) shows the optimality gap across iterations. We implemented the MAML-LQR on three scenarios: the full task pool (40 tasks), a selected subset (10 tasks), and two ablation baselines - selected subset without weight assignment and randomly selected subset. Our results demonstrate the faster convergence on the weighted selected subset, while both unweighted selected subset and random subset achieves at most the same performance as the full task pool. Moreover, Figure 3 (middle) depicts the optimality gap between the selected subset and full task pool with respect to sample size, confirming our theoretical results on sample complexity reduction for the MAML-LQR (Corollary 2). Figure 3 (right) shows the task-specific optimality gap on unseen meta-testing tasks drawn from the same training task distribution $\mathcal{P(T)}$. The results show that the learned meta-controller on $\mathcal{S}$ achieves comparable generalization performance while reducing sample complexity in the meta-training, which outperforms the randomly initialized controller.
@@
Figure 3: Optimality gap of Algorithm 1 in the MAML-LQR setting with respect to iterations and number of samples.
@@ Acknowledgements
Leonardo F. Toso is funded by the Center for AI and Responsible Financial Innovation (CAIRFI) Fellowship and by the Columbia Presidential Fellowship. James Anderson is partially funded by NSF grants ECCS 2144634 and 2231350, and the Columbia Data Science Institute.